OpenStudy (anonymous):

Find the angle between the given vectors to the nearest tenth of a degree. u = <-5, -4>, v = <-4, -3>

OpenStudy (anonymous):

take the dot product.

OpenStudy (anonymous):

so <20, 12>?

OpenStudy (anonymous):

if a and b are two vectors ,then \[vector a. vector b=\left| a \right|\left| b \right|\cos \theta \] where theta is the angle between the vectors.

OpenStudy (anonymous):

so how would I start that?

OpenStudy (zzr0ck3r):

@surjithayer `\(a\cdot b = ....\)`

OpenStudy (zzr0ck3r):

\(\overline{a}\cdot \overline{b}=||\overline{a}|| \ ||\overline{b}||\cos(\theta)\)

OpenStudy (zzr0ck3r):

:)

OpenStudy (anonymous):

u=-5 i-4 j v=-4 i-3 j \[u.v=\left| u \right|\left| v \right|\cos \theta \] \[u.v=(-5i-4j).(-4i-3j)=(-5)(-4)+(-4)(-3)=20+12=32\] \[\left| u \right|=\sqrt{\left( -5 \right)^2+\left( -4)^2 \right)}=\sqrt{25+16}=\sqrt{41}\] similarly find \[\left| v \right|\] and solve by the above formula.

OpenStudy (anonymous):

What answer did you get? because i got sqrt of 37

OpenStudy (anonymous):

I know it was wrong.